### v1

# BED-LLM: Intelligent Information Gathering with LLMs and Bayesian Experimental Design

Abstract
We propose a general-purpose approach for improving the ability of Large Language Models (LLMs) to intelligently and adaptively gather information from a user or other external source using the framework of sequential Bayesian experimental design (BED). This enables LLMs to act as effective multi-turn conversational agents and interactively interface with external environments. Our approach, which we call BED-LLM (Bayesian Experimental Design with Large Language Models), is based on iteratively choosing questions or queries that maximize the expected information gain (EIG) about the task of interest given the responses gathered previously. We show how this EIG can be formulated in a principled way using a probabilistic model derived from the LLM’s belief distribution and provide detailed insights into key decisions in its construction. Further key to the success of BED-LLM are a number of specific innovations, such as a carefully designed estimator for the EIG, not solely relying on in-context updates for conditioning on previous responses, and a targeted strategy for proposing candidate queries. We find that BED-LLM achieves substantial gains in performance across a wide range of tests based on the 20–questions game and using the LLM to actively infer user preferences, compared to direct prompting of the LLM and other adaptive design strategies.
1 Introduction
Despite their many successes, LLMs currently fall short on a crucial aspect of interactive intelligence: proactively seeking out information from a user or external environment in an intelligent and adaptive manner (Laban et al,, 2025; Li et al, 2025c, ). For example, they have been shown to perform poorly on problems such as multi-step guessing games (Bertolazzi et al,, 2023; Zhang et al,, 2024), task clarification (Chi et al,, 2024), IT task automation (Jha et al,, 2025), and iterative external tool use (Patil et al,, 2025). In particular, while modern LLMs are often capable of producing coherent and insightful questions (or other external queries) in a one-shot manner, they typically struggle to appropriately tailor their questions to previously gathered responses on interactive tasks (Bertolazzi et al,, 2023; Patil et al,, 2025).
There is, therefore, a pressing need to improve the ability of LLMs to adaptively ask questions and gather information in a targeted manner. Such capabilities are essential for a wide variety of problems, such as clarifying user intent, personalizing model behavior to a particular user, or generally acting as effective multi-turn conversational agents. They are also critical if we want to use LLMs in data gathering tasks or as automated agents in decision-making pipelines (Wu et al,, 2025). In turn, these capabilities are essential for a wide variety of downstream applications from medical diagnostic assistants, targeted code generation, and tutoring systems, to conducting automated surveys, AI–driven scientific inquiries, and market research (Liu et al, 2024a, ; Mandal et al,, 2025; Jacobsen et al,, 2025; Lu et al,, 2024). Note that in all these problems it is not enough for the LLM to simply one-shot generate good fixed questions based only on its parametric knowledge. Instead, we need it to be able to adaptively choose questions that are good, conditioned on the already-collected user responses.
We propose to address this challenge using the framework of sequential Bayesian experimental design (BED, Lindley, (1956); MacKay, (1992); Chaloner & Verdinelli, (1995); Sebastiani & Wynn, (2000); Rainforth et al, (2024)), which provides a model-based, information-theoretic mechanism for making adaptive design decisions, given a generative model of the experiment. Specifically, we show how the problem of interactive information gathering with LLMs can be formulated as a sequential experimental design problem, wherein we iterate between choosing queries based on maximizing their expected information gain (EIG) and updating our beliefs with the information from the received response. Here, the underlying generative model from which the sequential BED procedure is formed is derived from the LLM, and we carefully show how this should be constructed, providing extensive insights on key design decisions.
We call our approach BED-LLM (which stands for Bayesian Experimental Design with Large Language Models) and show how its success is critically dependent on careful choices for how the EIG is estimated, how candidate queries are generated, and how the model is updated given previous responses. For example, we show how approximations to the EIG made in previous work (Kobalczyk et al,, 2025; Hu et al,, 2024) can significantly degrade performance compared to using more carefully constructed estimation strategies. We also show how simple in-context learning (Brown et al,, 2020) is typically not sufficient to incorporate all the information from previous query–response pairs, with an additional sample–then–filter strategy introduced to correct this. Finally, we introduce a targeted strategy for proposing candidate queries (whose EIG we then evaluate) that directly utilizes our current belief state and provides important improvements over direct generation by the LLM.
Together, we find that these innovations provide substantial performance benefits over directly generating queries from the LLM and more basic approximations of the sequential BED framework. Specifically, we first find that BED-LLM provides substantial improvements in the success rate for the 20-Questions problem across a variety of LLMs and target quantities. For example, when predicting celebrities with Mistral-Large, we observe an improvement in success rate from 14% to 91%. Second, we demonstrate noticeable improvements in using the LLM for movie recommendations, showing that these benefits hold even when the LLM’s predictive model differs from that of the answerer.
In summary, we illustrate how the adaptive information gathering capabilities of LLMs can be significantly improved by the careful application of the sequential BED framework to their decision making, thereby making an important step towards their more general usage as interactive agents.
2 Problem Formulation and Background
There are two natural ways one might look to improve LLMs’ adaptive information gathering capabilities: we can look to change the LLM itself through finetuning or altering the pretraining process, or we can make changes in how the LLM is used at deployment time. Our focus will be on the latter, as this means we can make improvements to performance without requiring access to any additional data or expensive retraining, noting that for information–gathering problems, we will not usually have task–specific data available upfront (e.g. we cannot expect our LLM to have any knowledge about a particular user’s preferences). However, we emphasize that any improvements that can be made to the LLM itself will be complementary to our deployment–time approach.
To formalize the notion of information gathering, we need a concrete idea of what we wish to learn about. We denote this target quantity as $\theta$, and note that it could, for example, be a user’s preference, the answers to one or more questions, a recommendation, or a piece of content that a user wants to generate. To meaningfully reason about $\theta$, we need to have some belief distribution for it, which we denote $p\left(\theta;h_{t}\right)$, [^1] Note we carefully distinguish between explicit conditioning, i.e. $p(a|b)$, and more general dependency, $p(a;b)$, in our notation. The former explicitly corresponds to the conditional distribution of an associated joint distribution, $p(a,b)$, while the latter may not. So here, for example, $h_{t}$ influences our distribution on $\theta$, but we are not deriving this distribution through conditioning of some explicit joint distribution. where $h_{t}$ represents the history of queries and responses at time $t$. As more queries are made, this history will grow and our beliefs on $\theta$ will update accordingly. We note that $p\left(\theta;h_{t}\right)$ will typically not be defined explicitly. Even in the LLM setting, there is significant flexibility to how it is constructed, as we discuss in Sections 3.3 and 4.
2.1 Information Gathering with LLMs
Our focus is on how to use LLMs to gather information about our target quantity $\theta$. The precise set of problems this can relate to is quite broad and includes, to name but a few, task clarification (Zhang et al, 2025a, ), preference learning (Handa et al,, 2024; Chakraborty et al,, 2024; Ouyang et al,, 2022), diagnosis and troubleshooting (Hirosawa et al,, 2024; Jha et al,, 2025), automating external simulations or real-world experiments (Mandal et al,, 2025), conducting surveys or collecting feedback (Aher et al,, 2023; Lee et al,, 2024), and AI tutoring agents (Kestin et al,, 2025; Liu et al, 2024a, ).
What all of these have in common is that the LLM starts with incomplete information about the target $\theta$, but can refine its beliefs by making queries, $x\in\mathcal{X}$, to the user or some other external agent and receiving responses, $y\in\mathcal{Y}$, that are informative about $\theta$. In general, multiple such queries, $x_{1},\dots,x_{n}$, can be made. We can choose them adaptively in a sequential decision-making process where we iteratively choose each $x_{t+1}$ based on the collected history $h_{t}:=(x_{i},y_{i})_{i=1:t}$. For clarity of exposition, we focus on the case where the $x_{i}$ correspond to explicit questions asked to the user, but emphasize that the approach applies more broadly to other forms of external interaction by the LLM, such as retrieving documents or calling external functions.
The simplest and cheapest way to incorporate the interaction history into the LLM to aid our sequential decision making is to include it in the context (Brown et al,, 2020). Namely, if the LLM’s distribution over generated text, $z\in\mathcal{Z}$, is $p_{\mathrm{LLM}}(z;c)$, where $c$ is an initial prompt, then we can consider $p_{\mathrm{LLM}}(z;[c,h_{t}])$ as an updated LLM that uses the previous question–response pairs as in-context examples, where $c,h_{t}]$ represents some appropriate concatenation of the history into the context. From this, we can derive an updated belief distribution over $\theta$. Most simply, this can be done by using $p_{\mathrm{LLM}}(z;[c,h_{t}])$ to directly query about $\theta$ (e.g. if $\theta$ is some preference, we could prompt the LLM to predict this preference). However, as we show later, this approach often fails to appropriately or fully incorporate the information from $h_{t}$, leading to belief states that are inconsistent with past observations. We will introduce a more robust and sophisticated method for deriving $p(\theta;h_{t})$ from $p_{\mathrm{LLM}}(z;[c,h_{t}])$ in Section˜3.3.
In the interest of brevity, for the rest of the paper we will drop the initial prompt $c$ from our notation and also assume that an appropriate final prompt has been added to end of the context to predict/sample our quantity of interest. For example, we will use $p_{\mathrm{LLM}}(\theta;h_{t})$ to denote the distribution when our context is $c,h_{t}]$ followed by an appropriate prompt to sample $\theta$ as its next token(s).
2.2 Information–Theoretic Experimental Design
Our aim of maximizing the information gathered about a target quantity can be formalized using the framework of Bayesian experimental design (BED; Lindley,, 1956; 1972; Chaloner & Verdinelli,, 1995; Rainforth et al,, 2024), which provides a powerful and principled framework for data acquisition.
The core of the BED framework is a joint generative model, $p(\theta,y;x)$ over the target quantity $\theta$ and outcomes, $y$, given designs $x$. Traditionally, this is most commonly specified as a conventional Bayesian model using a prior $p\left(\theta\right)$ and likelihood $p(y|\theta;x)$ (Rainforth et al,, 2024), but it can also be directly defined using a stochastic simulator (Kleinegesse & Gutmann,, 2020; Rainforth,, 2017) or as the marginal of some larger model (e.g. because $\theta$ corresponds to something other than the model parameters (Bickford Smith et al,, 2023; MacKay,, 1992)).
In the most general case, designs are then chosen to maximize the expectation of some utility function $U(\theta,y,x)$ under this generative model; that is we take $x^{*}=\mathrm{argmax}_{x}\mathbb{E}_{p(\theta,y;x)}\left[U(\theta,y,x)\right]$. The most common choice is to take $U(\theta,y,x)=\log p(\theta,y;x)-\log p\left(\theta\right)\log p(y;x)$ where $p\left(\theta\right)$ and $p(y;x)$ are the marginal distributions on $\theta$ and $y$ implied by our joint model and we have assumed that our current beliefs on $\theta$ are independent of the design. This leads to an objective corresponding to the expected information gain (EIG) in $\theta$ (Lindley,, 1956; 1972),
$\mathrm{EIG}_{\theta}(x)$ $=\mathbb{E}_{p(\theta,y;x)}\left[\log\frac{p(\theta,y;x)}{p\left(\theta\right)p(y;x)}\right],$ (1)
which using Bayes’ rule can also be represented in the following equivalent forms
$\text{EIG}_{\theta}(x)$ $=\mathrm{H}[p\left(\theta\right)]-\mathbb{E}_{p(y;x)}[\mathrm{H}[p(\theta|y;x)]]$ (2)
$=\mathrm{H}[p(y;x)]-\mathbb{E}_{p\left(\theta\right)}[\mathrm{H}[p(y|\theta;x)]],$ (3)
where $\mathrm{H}$ denotes the Shannon entropy (i.e., $\mathrm{H}[p\left(\theta\right)]=-\mathbb{E}_{p\left(\theta\right)}[\log p\left(\theta\right)]$). We can thus equivalently think of this EIG as the mutual information between $\theta$ and $y$, the expected information gain over possible data simulated from our model (where the information gain is defined as the reduction in entropy from our prior on $\theta$ to the posterior), and the expected reduction in entropy over data from observing $\theta$ simulated from our prior (Sebastiani & Wynn,, 2000). By maximizing this objective with respect to $x$, we are thus maximizing the expected amount of information we gain in $\theta$, subject to the assumptions that our generative model can both accurately estimate the true data-generating process and encapsulate our beliefs about $\theta$.
Working with the EIG is highly suited to a sequential or adaptive design approach, wherein it is generally referred to as sequential BED or Bayesian Adaptive Design (Rainforth et al,, 2024). Because the EIG is only a function of our underlying model, when we update the model as new data becomes available, our EIG design objective will naturally update as well. In particular, if $p(\theta,y_{t+1};h_{t},x_{t+1})$ represents our generative model after already observing the history $h_{t}$, then we can iteratively optimize the incremental EIG (MacKay,, 1992; Cavagnaro et al,, 2010),
$\mathrm{EIG}_{\theta}(x_{t+1};h_{t})$ $=\mathbb{E}_{p(\theta,y_{t+1};h_{t},x_{t+1})}\left[\log\frac{p(\theta,y_{t+1};h_{t},x_{t+1})}{p\left(\theta;h_{t}\right)p(y_{t+1};h_{t},x_{t+1})}\right].$ (4)
Here our updated model, $p(\theta,y_{t+1};h_{t},x_{t+1})$, conventionally comes from a Bayesian update of the original model, noting that the Savage axioms dictate that a rational agent should update its beliefs in a Bayesian manner (Savage,, 1954). However, in many cases, this is not practical and other non-Bayesian updates are performed instead. In particular, in many Bayesian active learning problems, the update actually corresponds to retraining the model with an updated dataset (Gal et al,, 2017; Bickford Smith et al,, 2023).
It has recently been noted that this traditional sequential BED approach can sometimes be suboptimal in practice, as it only optimizes the EIG of the next observation, without planning ahead for the fact that design decisions taken at a given step can also influence the achievable EIGs from future steps (Foster,, 2021). A variety of policy–based BED approaches have subsequently been proposed to address this (Foster et al,, 2021; Ivanova et al,, 2021; Blau et al,, 2022; Huan & Marzouk,, 2016; Hedman et al,, 2025), while also removing the need to make model updates and conduct optimizations during the experiment itself. However, training these policies typically requires a very large number of simulations from the generative model, which makes them impractical for our setting, due to the expense of simulating full experimental rollouts from the LLM.
3 Sequential Bayesian Experimental Design for LLMs
The sequential BED framework described in Section˜2.2 requires two core components to be specified by the user: an initial joint model $p(\theta,y;x)$ over hypotheses $\theta$ and outcomes $y$, given chosen experiment(s) $x$; and a procedure to derive an updated model $p(\theta,y_{t+1};h_{t},x_{t+1})$ after observing an experimental history $h_{t}$. In the LLM setting, there is significant flexibility in these critical design decisions. In particular, there are various ways that one can derive a suitable joint distribution from the LLM and the ability to learn in context provides opportunities for update methods that go beyond standard Bayesian model updates.
Model Construction
A major challenge in the LLM setting is that unlike conventional probabilistic models, in general, $p_{\mathrm{LLM}}(\theta)\,p_{\mathrm{LLM}}(y;[\theta,x])\neq p_{\mathrm{LLM}}(y;x)\,p_{\mathrm{LLM}}(\theta;[x,y])$. That is, we induce a different joint distribution if we first sample $\theta$ then sample $y$ with $\theta$ in context (which we refer to as the prior–likelihood pairing), than if we first sample $y$ then sample $\theta$ with $y$ in context (which we refer to as the data–estimation pairing). Moreover, in each pairing we may wish to deviate from the LLM’s directly induced distribution on one or both components. The success of using sequential BED with LLMs turns out to be critically dependent on which pairing we choose and the deviations we make from LLM’s directly induced distribution.
We delay proper discussion of this complex issue until Section˜4, where we will see that the preferable setup can depend on the precise problem setting and, in particular, the relative complexity of spaces of $\theta$ and $y$. For now, we will focus on using the prior–likelihood pairing, as this transpires to generally be the advantageous setup when the space of $\theta$ is more complex than that of $y$, which we believe to be the more commonly occurring setting and the one we consider in our experiments. While we will generally use the LLM’s directly induced distribution for the likelihood, we allow the prior to deviate from this in a problem–specific manner. As such, our initial joint model will be $p(\theta,y;x)=p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])$, where the specification of the prior $p(\theta)$ can vary between problems.
Model Updating
Optimally updating the joint model in this setting requires incorporating new observations in a way that both fully captures the available information from new data and is computationally tractable. At one extreme, we could perform full Bayesian updates via approximate inference, as in classical sequential BED. This approach, however, demands a prohibitively large number of LLM evaluations to accurately approximate the posterior, and it does not exploit the power of the LLM as a probabilistic generative model, where autoregressive sequential rollouts often lead to more nuanced and diverse behavior than repeated static likelihood queries. The other obvious choice is to rely on simple in context updates, such that we use $p(\theta;h_{t})=p_{\mathrm{LLM}}(\theta;h_{t})$. However, as we show later, this fails to sufficiently capture the information provided by the new data and thus undermines the sequential BED approach.
As we will discuss in Section˜3.3, we therefore employ a strategy that is somewhere between the two, wherein we set up a rejection sampling procedure that draws samples from the in–context distribution, $p_{\mathrm{LLM}}(\theta;h_{t})$, but then filters out samples that are actually not compatible with previous observations. We refer to the resulting distribution as $p_{f}(\theta;h_{t})$.
In our experiments and subsequent exposition of our approach, we will not update the likelihood model as more data is gathered, that is, our likelihood in the sequential setting will be $p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ instead of $p_{\mathrm{LLM}}(y_{t+1};[h_{t},\theta,x_{t+1}])$. The main rationale of this choice is that for many problems our beliefs on $\theta$ capture all the required information to predict $y|x$, hence including the history is adding unnecessary context that could influence the LLM’s behavior in undesirable ways. However, it is important to note that $p_{\mathrm{LLM}}(y_{t+1};[h_{t},\theta,x_{t+1}])$ should be used instead for problems where $\theta$ will not capture all information from previous data, e.g. if $\theta$ is a binary value corresponding to whether we reject a null hypothesis, or is the answer to a particular other question of interest.
BED-LLM
We now introduce our specific principled algorithm for intelligent and adaptive information gathering in LLMs using sequential BED. We call our approach BED-LLM, which stands for Bayesian Experimental Design with Large Language Models. Here, the questions posed to the user [^2] More broadly, these could be more general queries such as document retrieval or function calling; however in our experiments we only consider questions asked of a user. will correspond to our designs, $x$, and the responses received will correspond to our outcomes, $y$. Using the LLM to derive a joint model over these outcomes and the target variables $\theta$ as described above, we can derive a precise form of the expected information gain, $\text{EIG}_{\theta}(x)$. We can then interleave choosing informative questions by optimising the incremental EIG defined in Equation˜4, and updating our underlying model based on the received question–response pairs. Specifically, BED-LLM iterates over the following key steps, where $t$ indexes the current turn.
  1. 1.
    Generate Candidate Questions: Propose a candidate set of $M$ diverse, multiple-choice questions, $\mathcal{X}^{\text{cand}}$, by appropriate sampling of the LLM based on the current conversational context (Section˜3.1).
  2. 2.
    Compute EIG Estimator: For each candidate question $x\in\mathcal{X}^{\text{cand}}$, estimate $\text{EIG}_{\theta}(x)$ as described in in Section˜3.2.
  3. 3.
    Select and Ask Optimal Question: Choose the question $x_{t}\in\mathcal{X}^{\text{cand}}$ that yields the highest estimated EIG. Pose $x_{t}$ to the user, observe response $y_{t}$, and update the history, $h_{t}=(h_{t-1},(x_{t},y_{t}))$.
  4. 4.
    Construct updated joint, $p(\theta,y_{t+1};h_{t},x_{t+1})=p_{f}(\theta;h_{t})p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$, using the new history (Section˜3.3) and return to Step 1 (unless a termination criterion like maximum number of questions has been achieved).
Because BED-LLM uses the prior–likelihood pairing, our belief state on $\theta$ is simply given by $p_{f}(\theta;h_{t})$ at any given point in the process; our underlying aim is that the uncertainty of this distribution contracts as quickly as possible.
In the remainder of this section, we provide details on these key individual steps. We note that some of the choices made are closely linked to our joint model construction; we will later provide extensive discussion on why this model construction was chosen and possible alternatives in Section˜4.
3.1 Candidate Question Generation
The first step of each BED-LLM iteration is to generate a finite pool of candidate questions, $\mathcal{X}^{\text{cand}}$, from which our chosen question will be selected. To ensure we ultimately choose good questions, we must generate a candidate set $\mathcal{X}^{\text{cand}}$ that is rich and diverse enough to include sufficiently informative questions in the first place.
We consider two approaches for generating candidate questions:
  • Unconstrained generation. Given the accumulated history, the LLM is simply asked to propose new questions by sampling from $p_{\mathrm{LLM}}(x_{t+1};h_{t})$ with appropriate prompting.
  • Conditional generation. The LLM is given both the accumulated history and a generated set of hypotheses $\Theta^{\text{cand}}=\{\theta^{(n)}\}_{n=1}^{N}$, such that we sample from $p_{\mathrm{LLM}}(x_{t+1};[h_{t},\Theta^{\text{cand}}])$. Specifically, the LLM is prompted to propose questions whose answers would “slice” the hypothesis pool into roughly balanced subsets.
For both strategies, we sample $M$ questions independently with a relatively high temperature (we use $T=1.3$ in our experiments), in order to encourage diversity. Conditional generation allows us to “guide” the LLM to propose highly informative questions. However, it risks over-fitting to the sampled set of hypotheses. In practice, we found that conditional generation works well for discrete, easily divisible hypothesis spaces such as those found in the 20–Questions setting, but tends to overfit in the case of more complex, overlapping hypotheses such as persona identification. We can avoid this overfitting by generating two hypothesis sets, one for generating questions and one for estimating EIG. However, this adds significant computational cost, leading us to prefer the unconstrained generation approach in such cases.
We restrict our questions to be multiple-choice questions. This is because having a small space of single-token responses $y$ (e.g., yes/no, A/B/C/D) ensures that $p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ provides a meaningful uncertainty over outcomes and that we can reliably calculate/estimate its entropy. We discuss this restriction at length in Section˜4.2, where we will see that limitations in LLM uncertainty quantification over unrestricted spaces generally require us to use a restricted space for $y$ (as done here) or $\theta$ (which can be tackled by instead using the data–estimation pairing).
3.2 Estimating EIG for each question
We can estimate the EIG for a given question $x_{t+1}$ using the following Rao-Blackwellized estimator of Equation 3, independently introduced by Gal et al, (2017) and Rainforth, (2017):
$\begin{split}\text{EIG}_{\theta}(x_{t+1};h_{t})\approx&\frac{1}{N}\sum_{n=1}^{N}\sum_{y_{t+1}\in\mathcal{Y}}p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])\log p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])\\ &-\sum_{y_{t+1}\in\mathcal{Y}}\hat{p}(y_{t+1};[h_{t},x_{t+1}])\log\hat{p}(y_{t+1};[h_{t},x_{t+1}]),\end{split}$ (5)
where $\hat{p}(y_{t+1};[h_{t},x_{t+1}])\!:=\!\frac{1}{N}\sum_{n=1}^{N}p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])$ and $\theta^{(n)}\sim p_{f}(\theta;h_{t})$ (see Section˜3.3). While this estimator is biased for finite NN, [^3] In general it is not possible to derive unbiased estimates for the EIG without using MLMC approaches (Goda et al,, 2022) that significantly increase the cost-normalized variance. it is still a singly intractable estimator that converges as $N\to\infty$ at the standard Monte Carlo rate $O(1/\sqrt{N})$ (Rainforth et al,, 2018). Note that the samples $\theta^{(n)}$ do not need to be independent for this to hold, provided that they can be shown to satisfy some appropriate form of ergodicity or decaying correlation (see, e.g., Billingsley, (2013)).
Note that when constructing this estimator, one should directly utilize the LLM’s logits (whenever this is possible) to calculate the $p_{\mathrm{LLM}}(y_{t+1};[\theta^{(n)},x_{t+1}])$ terms. By the Rao-Blackwell theorem (Rao et al,, 1945), this will always produce lower variance estimators than purely sample–based estimators, like those employed in Hu et al, (2024) and Kobalczyk et al, (2025).
3.2.1 EIG cannot be well approximated using predictive entropy
Previous attempts to apply BED to the LLM setting with a prior–likelihood based model have all, either explicitly or implicitly, assumed that the expected likelihood entropy, $\mathbb{E}_{p(\theta;h_{t})}[\mathrm{H}[p(y_{t+1}|\theta;x_{t+1},h_{t})]]$, is constant over designs (Kobalczyk et al,, 2025; Mazzaccara et al,, 2024; Cooper et al,, 2025); for BED-LLM this would equate to ignoring the first line of the estimator in Equation˜5. Under this assumption, the EIG breaks down to simply the marginal predictive entropy, $\mathrm{H}[\mathbb{E}_{p(\theta;h_{t})}[p(y_{t+1}|\theta;x_{t+1},h_{t})]]$. Therefore, approaches that make this assumption are no longer information–based, and are instead making decisions simply based on predictive uncertainty.
This is problematic as, in practice, this expected conditional uncertainty will vary between questions. For example, if $\theta=$“dog”, then the response to the question “Is it an animal?” should be close to deterministic, but the expected response to the question “Does it have black fur?” clearly is not. In general, $\mathbb{E}_{p(\theta;h_{t})}[\mathrm{H}[p(y_{t+1}|\theta;x_{t+1},h_{t})]]$ measures how easily the question can be answered once $\theta$ is known, and intuitively this will not be constant across questions for the vast majority of problems.
This variability in the expected entropy of the likelihood can be crucial in determining good questions and omitting it provides no meaningful computational saving (as the same LLM evaluations are used for the top and bottom lines of Equation˜5, doing the full estimate of the EIG requires no additional LLM calls to be made). We, therefore, strongly advocate against making this assumption: including the expected entropy of the likelihood is essential in avoiding questions that are irrelevant, ambiguous, unclear, or simply unhelpful in our quest to learn about $\theta$. This recommendation will be strongly backed up by our experiments, where we will see that approximating the EIG with the marginal predictive entropy significantly degrades performance of the sequential BED approach. A further example of how predictive entropy can be detrimental is given in Figure˜1.
One might instead be tempted to replace $\hat{p}(y_{t+1};[h_{t},x_{t+1}])$ with $p_{\mathrm{LLM}}(y_{t+1};[h_{t},x_{t+1}])$ in the EIG estimation, as the two essentially offer alternative predictive distributions for the outcome. We also advise against this though, noting that it again provides no meaningful computational benefits (unless one also assumes a deterministic likelihood, but this would then mean we no longer consider $\theta$ at all). A key reason for avoiding this substitution is that it would mean we are no longer estimating a true EIG: the inconsistency between the likelihood and the marginal data distribution means there is no longer a joint model where we are minimising our expected uncertainty in $\theta$. We also find that the LLM process of sampling $\theta$ from $p_{f}(\theta;h_{t})$ followed by $y$ from $p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ tends to give a better uncertainty over responses than sampling $y$ directly from $p_{\mathrm{LLM}}(y_{t+1};[h_{t},x_{t+1}])$.
3.3 Prior Construction and Belief Updating
The Savage axioms (Savage,, 1954) tell us that a rational agent should update its beliefs in a Bayesian manner. However, doing full Bayesian updates to our model as the history grows is generally impractical for computational reasons in the LLM setting, as it requires approximate inference and this, in turn, typically requires large numbers of expensive likelihood evaluations. Furthermore, the Savage axioms only hold if our (implied) prior truly represents our beliefs, but we find that $p_{\mathrm{LLM}}(\theta)$ is typically heavily over–collapsed to a small number of possible hypotheses and can struggle to convey the full range of possibilities even with careful prompting and a high temperature (c.f. Figure˜8).
Question 1 Question:
Which ice cream flavor feels like the best match for this user?
Choose one option: A. Vanilla B. Dark Chocolate C. Strawberry Swirl D. Mint Chocolate Chip
Predictive Entropy: Very High
EIG: 0
Question 2 Question:
Which film genre does the user most prefer?
Choose one option: A. Action B. Sci-Fi C. Comedy D. Horror
Predictive Entropy: High
EIG: High
Figure 1: Predictive entropy vs. expected information gain (EIG) in a film‐preferences elicitation task. Left: very high predictive entropy (answer is completely unknown) but EIG =0=0 because the answer provides no insight into the user’s film preferences. Right: both predictive entropy and EIG are high as the answer is uncertain, but different answers would lead to markedly different posterior updates, making it informative for learning film preferences. This thus demonstrates how the two criteria can select different questions.
A natural tractable alternative is to simply derive our beliefs through in-context updates, that is, use $p_{\mathrm{LLM}}(\theta;h_{t})$, noting that this has been shown to behave differently to Bayesian updating in the LLM context (Falck et al,, 2024; Kossen et al,, 2024). Unfortunately, we find that the ability of current LLMs to fully incorporate the information from $h_{t}$, even when using large state-of-the-art LLM variants such as GPT-4o (OpenAI,, 2024) or Mistral Large (Mistral AI Team,, 2024), is surprisingly lacking. In particular, we find that all LLMs we evaluate regularly sample hypotheses $\theta$ that are incompatible with previous observations in $h_{t}$, especially as the length of the history grows. We also, again, find problems with over–collapsing, with LLMs often jumping to conclusions about the value of $\theta$ without appropriate evidence for this certainty in $h_{t}$.
The shortfalls of in-context learning in such settings have also previously be noted by, for example, (Liu et al, 2024b, ; Zhang et al, 2025b, ; Zhang et al,, 2024). We posit two reasons why they likely struggle in such settings. First, the information from the different examples in the history are generally highly distinct in these information-gathering settings (indeed, this is part of our aim in adaptively design informative questions), making it harder for the LLM to appropriately reconcile all the provided information than in many other uses of in–context learning. Second, $\theta$ will often represent a user–specific variable that cannot easily be predicted from any data other than the user’s responses to questions: it has been argued that much of the success of in–context learning in LLMs is down to improving problem specification and linking the requested task to data it has seen in its training, rather than truly “learning” from the provided examples (Min et al,, 2022; Kossen et al,, 2024), but the history in our setting is rarely helpful for this due to its user–specific nature.
Addressing these shortfalls is critical to the success of BED-LLM. Although we will still use $p_{\mathrm{LLM}}(\theta;h_{t})$ as the basis for deriving our belief state over $\theta$ (i.e. our intermediate prior), we make various alterations to effectively incorporate historical information and ensure diversity. We refer to our derived distribution as $p_{f}(\theta;h_{t})$ and its key component is to filter the generated hypotheses according to whether they are compatible with the history $h_{t}$. We do this by using the LLM to zero-shot check the compatibility of each sampled $\theta$ with all the previous question–answer pairs in $h_{t}$ (using $p_{\mathrm{LLM}}(y;[\theta,x])$) and then rejecting that sample if an incompatibility is found. Specifically, a sample is rejected if the likelihood of an observed answer falls below a predefined threshold, chosen to balance robustness to model uncertainty against the need to enforce strict historical coherence. To reduce the computational cost of generating and evaluating hypotheses, we further include a hypothesis-retention mechanism: any hypotheses from the previous step t1t\!-\!1 which remain consistent with the most recent question and observation are retained in the hypothesis set for step $t$ without regeneration.
We also make some further variations from $p_{\mathrm{LLM}}(\theta;h_{t})$ to try and combat the over–collapsing issue and encourage diversity (noting that our initial sampling acts like a proposal in a rejection sampler and invalid samples should still be removed by the filtering). First, we sample at a high temperature ($T=1.3$). Second, rather than generating candidates independently, we prompt the LLM to generate batches of candidates using a prompt that requests a “diverse and representative set of candidates”. We can sample multiple batches, with the previous batches in context to force the LLM to generate different samples. Finally, we assume a uniform distribution over the $\theta$ that make it through the rejection, such that we remove any duplicated samples and do not make any direct use of the LLM’s logits on $\theta$. We note that the precise prompting strategy used in $p_{\mathrm{LLM}}(\theta;h_{t})$ can also be quite important in achieving effective performance, with our prompts set up to try and maximize the diversity in the generated $\theta$ samples. Details on our exact setup for $p_{f}(\theta;h_{t})$ are in Appendix˜B.
Under the aforementioned setup for $p_{f}(\theta;h_{t})$, the original prior $p(\theta)$ only affects the EIG at the first step as it has no direct bearing on $p_{f}(\theta;h_{t})$ for $t>0$. We, therefore, define our original prior to simply be equal to $p_{f}(\theta;h_{t})$ with $h_{t}=\emptyset$, such that no filtering is performed. We note though that in some cases it may be preferable to specify some customized $p(\theta)$, potentially then adjusting the filtering process and/or weighting the accepted samples in a non–uniform manner to ensure this prior information is not lost as more data is gathered.
4 On the Specification of $p(\theta,y_{t+1};h_{t},x_{t+1})$, and its Implications
As we described in Section˜3, adapting BED to the LLM setting hinges upon how we specify, and update, the joint distribution $p(\theta,y_{t+1};h_{t},x_{t+1})$. As we discuss in this section, this choice is critical to the success of BED-LLM and impacts how we model the current uncertainty about $\theta$; how easily and accurately we can estimate EIG; and how aligned the model’s actual information gain under a question is predicted by the EIG, which ultimately governs whether the questions we choose are actually good at reducing uncertainty about $\theta$.
As discussed at the start of Section˜3, there are two distinct ways to derive the joint model from our LLM: using a prior–likelihood pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p(\theta;h_{t})p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$}, or a data–estimation pairing, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$}, [^4] One could in principle also consider prompting the LLM to generate $\theta$ and $y$ jointly as well, but we will omit these from our discussions on the basis that this will generally be impractical from an estimation perspective, while LLMs’ autoregressive nature means that this should be equivalent to one of the aforementioned joints, up to effects in prompting strategy and API handling etc. where each of the $p$ terms either corresponds directly to the LLM’s distribution over outputs given the provided context and suitable prompting (i.e. the corresponding $p_{\mathrm{LLM}}(\cdot;\cdot)$), some valid distribution derived from this, or an externally specified distribution. The first construction mirrors deriving our beliefs about $\theta$ from a conventional Bayesian posterior with a concrete prior and likelihood derived (at least partially) from the LLM, whereas the second is analogous to a one-sample martingale posterior that directly samples hypothetical data and draws inferences on $\theta$ given hypothetical data using in-context learning (Fong et al,, 2023; Falck et al,, 2024).
In our outlined BED-LLM approach (c.f. Section˜3), we adopted a prior–likelihood pairing. Below, we will expand on some reasons to prefer this formulation and discuss certain settings where the data–estimation might be preferable.
4.1 Modeling flexibility
The most obvious relative merits of the prior–likelihood and data–estimation pairings are in the flexibility in how each term is chosen. The prior–likelihood pairing gives us greater flexibility to construct a prior set of beliefs over $\theta$ that is distinct to the LLM’s internal beliefs, as it allows us to directly control this prior by changing $p(\theta;h_{t})$, whereas the prior is only implicitly defined in the data–estimation pairing. In Section˜3.3 we exploited this flexibility through our definition of $p_{f}(\theta;h_{t})$, wherein we made various adjustments to $p_{\mathrm{LLM}}(\theta;h_{t})$ to encourage diversity and better incorporate information from the history.
On the other hand, the data–estimation pairing could provide some beneficial flexibility in specifying how the data itself is simulated through changing $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$. Most notably, in some applications like scientific discovery, we may have access to external simulators that can be run to predict our data, with the LLM used instead primarily to automate how we draw conclusions from such hypothetical data (Ghareeb et al,, 2025).
In the cases considered in this paper, we do not have such external, validated simulators, and the flexibility in prior construction is more important for our needs.
4.2 Faithfulness of conditional distributions
In Section˜4.1, we considered flexibility in specifying the marginal distributions $p(\theta;h_{t})$ and $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$. While deviations from relying on direct LLM predictions are also in principle possible for the conditional models $p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ and $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$, in practice, these will typically be more difficult and expensive to implement. This is first because these conditionals need to be instantiated for each sampled instance of the conditional variable ($\theta$ and $y_{t+1}$ respectively), rather than just needing us to set up one marginal distribution that can then be repeatedly sampled from. Second, to construct estimators for Equations (2) and (3), we require access to concrete probabilities for the conditional distributions (in order to calculate entropies), whereas we only needed to draw samples for the marginal distributions (in order to approximate expectations). As such, the conditionals need to be explicit distributions, or at least distributions where the probability can be cheaply estimated (noting that such estimation will be nested and thus more expensive than in conventional settings (Rainforth et al,, 2018)), so they are more difficult to define through the output of some algorithmic procedure, especially if the number of possible realizations is large.
When considering the conditional distributions, the decisive question on the relative merit of the two formulations is which conditional factor we are willing to trust the LLM to supply as a full probability distribution. Critically, we rely on how the LLM captures uncertainty in this full distribution—including, for example, tail behavior—not merely the fidelity of typical samples; the marginal factors, by contrast, only need to be sampled from. If we accept the LLM’s direct predictive distribution for $p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$, then we are basing our notion of uncertainty around (and will need to calculate) $\mathrm{H}[p_{\mathrm{LLM}}\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)]$, and if instead we place more faith in the LLM’s internal distribution for $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$, then we are basing our uncertainty around $\mathrm{H}[p_{\mathrm{LLM}}(\theta;[h_{t},x_{t+1},y_{t+1}])]$. In essence, the choice between prior–likelihood and data–estimation pairings thus comes down to whether we believe the LLM will produce a more appropriate conditional uncertainty over $\theta$ or $y$, along with our ability to numerically estimate this uncertainty cheaply (noting the nesting means we need to repeatedly carry out this estimation for different $\theta$ and $y$ samples respectively).
This difference becomes particularly noticeable when the complexities of the spaces of $\theta$ and $y$ differ significantly. Our ability to draw sensible samples of either will generally be quite robust to these spaces being complex or high–dimensional; this is where LLMs tend to thrive, effectively generating highly complex outputs in an autoregressive manner. However, evaluating the entropy of a distribution becomes dramatically harder as the dimensionality or complexity increases (Acharya et al,, 2019; Paninski,, 2003), and the entropy of the predictive distribution of an LLM in such cases will not typically provide a sensible measure of uncertainty (Kadavath et al,, 2022; Desai & Durrett,, 2020). As such, the decision between joint formulations should predominantly be based on the complexity of the space of $\theta$ versus that of $y$: we should generally favor the prior–likelihood formulation if $\theta$ is more complex and the data–estimation formulation if $y$ is more complex. For example, suppose $\theta$ represents the space of Python code a user wants to generate and $y$ is the answer to a simple yes/no question about that code’s desired behavior. Here, the LLM’s entropy over yes vs. no when given the code will be a much more meaningful uncertainty measure than its entropy over code when told the answer to the question.
For the problems that we consider, the space of $y$ is less complex than that of $\theta$, indicating we should, in general, use the prior–likelihood formulation. However, in cases where this is not true, the data–estimation formulation may be preferable instead. We provide additional discussion on the impact of the chosen pairing on our entropy estimate, plus discussion on the choice of $\theta$, in Appendix˜A.
4.3 Extracting the belief state
A further advantage of the prior–likelihood construction is that our belief state on $\theta$ can be extracted directly, as it is simply $p_{f}(\theta;h_{t})$. With the data–estimation construction, we would have to estimate a marginal distribution from $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$} by integrating out the synthetic response $p(\theta;h_{t})=\int{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)}\;\mathrm{d}y_{t+1}$. Direct access to $p(\theta;h_{t})$ is also important to ensure that our current belief state is independent of the next question $x_{t+1}$, as the marginal obtained from the integral above is unlikely to remain the same for all $x_{t+1}$ in the way we intuitively want it to.
4.4 Alignment between EIG and belief updating procedure
Our ultimate goal is to minimize uncertainty in $\theta$, as measured by its entropy. With this in mind, we can use the expected uncertainty reduction framework of Bickford Smith et al, (2025) to provide insights into how well our EIG formulation and belief updating procedures align.
To simplify discussions, for now we consider the setting where we choose a single question $x$ and obtain a response $y$. Following Bickford Smith et al, (2025), we can think of the “true” optimal design as selecting
$x^{*}_{\mathrm{true}}=\mathrm{argmin}_{x}\mathbb{E}_{p_{\mathrm{true}}(y;x)}\left[\mathrm{H}[p(\theta;x,y)]\right],$ (6)
where $p_{\mathrm{true}}(y;x)$ is the true response distribution and $p(\theta;x,y)$ is our belief state after the experiment.
Note here that true optimal design has no direct dependency on our current beliefs about $\theta$; it only depends on $p_{\mathrm{true}}(y;x)$ and the hypothetical beliefs we produce for given observed data, $p(\theta;x,y)$. Thus, we can now see that our choice of joint model corresponds to different choices for approximating these quantities. Assuming that the LLM distribution is used directly for the conditional as per Section˜4.2, we thus have that:
  • The prior-likelihood pairing equates to the approximations $p_{\mathrm{true}}(y;x)\approx\int p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])d\theta$ and $p(\theta;x,y)\approx p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])/\int p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])d\theta$;
  • The data–estimation pairing equates to directly specifying a model for $p_{\mathrm{true}}(y;x)$ and then using the approximation $p(\theta;x,y)\approx p_{\mathrm{LLM}}(\theta;[x,y])$.
The appropriateness of each of these options, therefore, comes down to how faithful these approximations are respectively to the true data distribution, $p_{\mathrm{true}}(y;x)$, and how we actually derive our belief distribution on $\theta$ in practice once we have seen the new data.
The former of these considerations is difficult to control for as we simply do not know the true response distribution and it is hard to say which approach will thus estimate it best (though we can refer to the discussion in Section˜4.1 to determine which best matches our beliefs about the true response distribution). However, we do know upfront how we plan to derive our belief distribution on $\theta$ in practice, so we can use this to guide which joint model we formulate our EIG from. Namely, we see that using
  • The prior–likelihood EIG pairing equates to assuming we will make a Bayesian update to our beliefs on $\theta$ using the likelihood $p_{\mathrm{LLM}}(y;[\theta,x])$;
  • The data–estimation EIG pairing equates to assuming we will make an in-context update to our beliefs on $\theta$, as we are treating $p(\theta;x,y)$ as $p_{\mathrm{LLM}}(\theta;[x,y])$.
Our preference between the pairings should therefore be guided in part by how we plan to update the model in practice. In particular, if we plan to make pure Bayesian updates, then the prior-likelihood formulation will tend to yield an EIG that is more faithful to our updating procedure, while if we only make simple in-context updates, the data–estimation formulation will tend to yield a more faithful EIG instead.
The update we use in practice, namely taking $p(\theta;x,y)=p_{f}(\theta;[x,y])$ as outlined in Section˜3.3, can be seen as being somewhere between the in-context and Bayesian updating: we initially sample from $p_{\mathrm{LLM}}(\theta;[x,y])$, but then perform filtering and other steps. The relative extent to which it resembles each will be problem–dependent and again be linked to how much we trust the LLM to capture uncertainty in the space of $\theta$ vs. $y$.
For the settings we consider, we expect $p_{f}(\theta;[x,y])$ to generally be better approximated by a Bayesian update than an in–context update, aligning with our decision to use the prior–likelihood formulation. The reasons for this are that a) the filtering often removes a large proportion of the generated samples, especially at later experiment turns, with $p_{\mathrm{LLM}}(\theta;h_{t})$ not fully incorporating information from the history; b) the maintaining of the set of one consistent hypotheses from one turn to the next encourages a more Bayesian behavior, with samples persisting unless contradicted by a new likelihood term; and c) the typical premature over–collapse of $p_{\mathrm{LLM}}(\theta;h_{t})$ to a small number of hypotheses means it is typically unrepresentative of our beliefs.
These theoretical benefits are perhaps secondary to the more practical benefits from the ease of constructing an appropriate model in the prior–likelihood formulation and avoiding direct uncertainty estimation in the space of $\theta$. Nonetheless, they help confirm that our choices have not induced unnecessary mismatch between the EIG formulation and our updating procedure.
The analysis here can get a little more complicated when we consider the fact that “one-step optimal” decisions like Equation˜6 may no longer be optimal in the sequential decision–making setting. We provide more discussion on this in Section˜A.3.
5 Related Work
In this work, we consider the broad problem of sequential information gathering using LLMs—that is, adaptively asking natural-language questions to allow the LLM to rapidly learn about a parameter of interest $\theta$. Several works have explored the baseline ability of LLMs to ask such questions (Zhang et al,, 2024; Li et al, 2025b, )—effectively, the “naive" baseline in our experiments. While these works demonstrate some ability to adaptively construct information-seeking questions, they often fail to extract important information (Li et al, 2025a, ), leading us to more sophisticated approaches.
Ours is also not the first paper to use a targeted model-based approach to information gathering with LLMs: several existing works have previously attempted to choose questions based on estimates of the expected information gain, marginal predictive entropy, or similar quantities (Piriyakulkij et al,, 2023; Hu et al,, 2024; Kobalczyk et al,, 2025; Wang et al,, 2025; Mazzaccara et al,, 2024; Chan et al,, 2025; Cooper et al,, 2025). However, as explained below, these existing methods all make design choices that significantly limit performance. Further, none of them provide comparable insights into how the joint model should be best formulated and updated from the underlying LLM; most of them even fail to directly link their approach and BED itself.
A number of these previous works, including all that also employ a prior–likelihood pairing in their joint model, actually implicitly select queries using predictive entropy in practice (rather than EIG as claimed) because they assume a constant expected likelihood entropy (Kobalczyk et al,, 2025; Mazzaccara et al,, 2024; Cooper et al,, 2025). As we discussed in Section˜3.2.1 and show empirically in Section˜6, this simplification is not generally reasonable—as predictive entropy is not in general a good approximation to EIG—and leads to worse performance. We believe our work to be the first which uses a prior–likelihood model pairing without simplifying the EIG to just the predictive entropy.
All of the other methods have instead relied on models derived from a data–estimation pairing, then either obtaining a sample-based estimate of $\mathbb{E}_{p\left(y_{t+1};[h_{t},x_{t+1}]\right)}\left[\mathrm{H}[p_{\mathrm{LLM}}\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)]\right]$ (Piriyakulkij et al,, 2023; Hu et al,, 2024; Wang et al,, 2025), or in the case of Chan et al, (2025), an approximation to this entropy based on the size of conformal sets. As discussed in Section˜4, our prior–likelihood model pairing provides significant benefits, particularly in the large or unbounded hypothesis spaces we explore in this paper. To get an approach which is based on the data–estimation model pairing to work, previous works have thus had to restrict the set of allowable hypotheses (i.e. $\theta$) to a small set, typically also restricting the space of allowing outcomes and further making a deterministic likelihood assumption (thereby again implicitly reducing the EIG to predictive entropy). These restrictions significantly limit applicability and performance.
Rather than augmenting a frozen LLM with the ability to estimate utility functions, some works have instead aimed to post-train an LLM to improve its ability to ask questions (Zhang et al,, 2024; Wu et al,, 2025; Andukuri et al,, 2024). Most do not explicitly consider informativeness of questions: Zhang et al, (2024) and Wu et al, (2025) use reinforcement learning techniques to reward generations that quickly lead to the correct answer, and Andukuri et al, (2024) builds on Li et al, 2025b by fine-tuning on successful traces. Mazzaccara et al, (2024) do indirectly incorporate uncertainty: they use predictive entropy to identify informative questions, and then either fine-tune on the highest-entropy question, or perform DPO comparing the highest-entropy question with a lower-entropy question. We do not address fine-tuning in this work, focusing instead on exploring the correct way to formulate BED using LLMs.
As discussed in Section˜4, a key challenge in adapting BED to the LLM setting is in aligning the expected information gain with the actual uncertainties extracted from the LLM after updating. Handa et al, (2024) take a different approach to this problem by using the LLM to generate features for an external conventional Bayesian joint model (in their case, a linear Bradley Terry model), rather than deriving their joint model more directly from the LLM itself. This can be a good choice when we already have a well-specified Bayesian model form for the problem at hand; however, this may be challenging in arbitrarily large and complex hypothesis spaces. In particular, their specific method is not applicable more widely beyond the preference learning context they consider.
6 Experiments
We now empirically evaluate our proposed BED-LLM method, compare it against relevant baselines, and analyze its properties through targeted ablation studies. We consider two interactive question–answering scenarios that test different types of sequential information acquisition: 1) 20–Questions: a deductive reasoning game requiring identification of a hidden target entity through binary (Yes/No) queries, and 2) Active Preference Elicitation: a task in which the goal is to infer a latent user preference vector via multiple–choice questions.
To evaluate the different questioner methods (i.e., BED-LLM and our baselines), we further need some form of answerer that represents the user or external environment we are trying to learn about. Due to the complications of evaluating with real human participants, we instead conduct our tests with a simulated user based on a separate LLM instance that has been given a ground truth target $\theta^{*}$ (corresponding to the hidden target entity for 20–Questions and a user profile for the active preference elicitation). Note that this answerer LLM does not itself perform any learning—it does not keep previous question–response pairs in context. While we consider both deriving the questioner and answerer from the same LLM and from different LLMs, we emphasize that they are always using separate LLM instances with different prompts and contexts, thus no information is transferred between them except in the questions and responses themselves.
We compare BED-LLM against two baselines:
a) Entropy: selecting questions using the maximum predictive entropy, as done in Kobalczyk et al, (2025), Mazzaccara et al, (2024), and Cooper et al, (2025). For fairness and to ensure a strong baseline, we keep the entropy setup as close to BED-LLM as possible (including prompts, candidate question generation, and our joint model construction with hypothesis filtering), varying only in replacing the EIG estimator with an estimator of the marginal predictive entropy,
$\mathrm{Entropy}_{\theta}(x_{t+1})=-\sum_{y_{t+1}\in\mathcal{Y}}\hat{p}(y_{t+1};[h_{t},x_{t+1}])\log\hat{p}(y_{t+1};[h_{t},x_{t+1}]).$ (7)
We construct an estimator for Equation˜7 analogously to Equation˜5. We also note here that our version of the Entropy baseline produces noticeably stronger performance on the 20–Questions problem than Kobalczyk et al, (2025) report for their own “EIG” approach, demonstrating it efficacy. [^5] Though Kobalczyk et al, (2025) describe their approach as targeting the EIG, the estimator they use (given by their Algorithm 1) actually equates to a sample-based estimator of the marginal predictive entropy instead and is thus comparable with our Entropy baseline. As they do not provide code specifically for the 20–questions experiments or explicit results on their approaches’ success rate, we were unable to provide direct comparisons against their exact setup. However, we believe that it is still clear from the results they do present that their approach is performing worse than the Entropy baseline we do consider. First, the reported ranks they give indicate they are not beating the naive LLM approach with the same regularity as Entropy does. Second, in their rebuttal on OpenReview they explain that they only report ranks and not success rate because “only very few runs would result in a non-zero accuracy”, whereas we see that our Entropy baseline is regularly successful.
b) Naive QA:we prompt the questioner LLM to directly produce an informative next question without explicit hypothesis generation, EIG computation, or filtering. The question is obtained by drawing a single sample at temperature $T=1$ from the questioner LLM. This approach, which was explored in Zhang et al, (2024), relies on the implicit reasoning capabilities of the LLM to select good questions.
6.1 20–Questions
We first evaluate BED-LLM in the classic “20–Questions” game, where the questioner must identify a hidden target $\theta^{*}$ by asking up to 20 Yes/No questions. Here, the hypothesis space of possible $\theta$ is large and not explicitly defined or restricted: it is bounded only by what the LLM can generate. As such, all previous work that has considered using BED-style decision with LLMs either a) cannot be directly applied in our precise problem setup because it requires the questioner to be given access to an explicit, restricted set of possible $\theta$ (Hu et al,, 2024; Piriyakulkij et al,, 2023; Wang et al,, 2025; Chan et al,, 2025), or b) is a variant of our Entropy baseline, due to using marginal predictive entropy in lieu of the EIG (Kobalczyk et al,, 2025; Mazzaccara et al,, 2024; Cooper et al,, 2025).
We evaluate across three distinct datasets each containing a mix of 100 obscure and common targets corresponding to either Animals, Celebrities, or Things. Here, the dataset is just a list of different $\theta^{*}$ that will be individually provided to the answerer to instantiate different problems (e.g. we conduct a trial where $\theta^{*}=$“dog”, then one where $\theta^{*}=$“cat”, etc). The list of targets is never provided to the questioner model to restrict the set of possible hypotheses: the questioner is only prompted that is trying to identify an “animal”, “celebrity”, or “thing”. The datasets are: a) Animals: a set of animal species generated with OpenAI’s o3 model to ensure a diverse mix and balanced taxonomy, see Appendix˜D for details; b) Celebrities: a diverse set of public figures, as used by Zhang et al, (2024); c) Things: a collection of everyday and exotic entities drawn from the web corpus, as used by Zhang et al, (2024); it covers a wide range of categories, from plants and clothing to professions, events, and mythical creatures. We note that the best success rates achieved by Zhang et al, (2024) on Celebrities and Things are substantially below those of both BED-LLM and our Entropy baseline.
Using our sample–then–filter process (c.f. Section˜3.3), we aim to sample at least $N=15$ hypotheses, repeating the cycle up to three times if needed (the exact number of hypotheses can be less than this as it may not be possible to generate sufficient valid hypotheses, especially in later experiment turns). The questioner generates $M=15$ candidate questions to test, $\mathcal{X}^{\text{cand}}$, using the “conditional generation” approach of Section˜3.1 when possible, but falling back on “unconditional generation” if insufficient candidate hypotheses have been generated.
Evaluation
[…]
Table 1: Success rate (%) for 20–Questions at the end of the game. Best result in bold. Success rates are integers because each dataset contains exactly 100 entities. $\pm$ numbers represent the standard error of the mean estimate (this is estimated using $\sqrt{p(1-p)/(n-1)}$ where $p$ is the success percentage and $n$ is the number of datapoints, noting this estimator is positively biased and thus conservative).
Success Rate (%) $\uparrow$
Dataset Model Naive QA Entropy BED-LLM
[…]
Figure 2: Success rate against turn for the 20–Questions game. Error bars represent the standard error of the mean estimate across 100 target entities in each dataset. The same model used as both questioner and answerer.
Figure 3: Success rates for the 20–Questions game with different models: the questioner Q and answerer A.
Results
First, we evaluate in the setting where the same LLM is used as both the questioner and the answerer. Table˜1 shows the success rates for the different datasets and a variety of different LLMs. Across all datasets and all LLMs, BED-LLM achieves the highest success rate, significantly outperforming the Entropy and Naive QA baselines. It is particularly notable how the success rate of our principled information-gain strategy is typically more than double that of Naive QA which relies solely on implicit LLM reasoning.
Per-turn success rates are shown in Figure˜2. Over the first few turns, BED-LLM and Entropy perform similarly. However, after around 7-10 turns we consistently see BED-LLM pull away from Entropy. This may simply be because it takes some time for the effect of choosing good questions to take effect (the large space of possible $\theta$ means that no approach can reliably determine the answer in only a few steps), or it may be that the effect of using proper EIG estimates becomes more pronounced as the set of viable hypotheses decreases and it is harder to find useful questions.
Different questioner and answerer
In the experiments above, the same LLM served as both the questioner and the answerer. This setting implicitly assumes that the questioner model has access to the answerer’s response distribution, i.e. that there is no misspecification in the likelihood used by BED-LLM and Entropy or the internal reasoning of Naive QA. However, in practice, the answerer is typically a human whose responses will follow a different distribution to the questioner’s internal model for reasoning about responses. To test the robustness of the various methods to this model misspecification, we therefore also evaluate scenarios where the questioner and answerer are formulated with entirely different LLMs, instead of just different instances of the same LLM. As shown in Figure˜3, the benefit of BED-LLM persists even under this model misspecification; it consistently achieves higher success rates than both baselines across datasets. In contrast, Entropy exhibits a more severe drop in performance in this more realistic setting, often now performing more in line with Naive QA, which itself appears to be quite robust to the misspecification.
Figure 4: Data–estimation ablation using Qwen2.5-72B-Instruct on Celebrities dataset.
Data–estimation ablation
As described in Section˜3, an important design choice is whether we choose the prior–likelihood pairing or the data–estimation pairing for BED-LLM. As an ablation for this critical decision, we also implement a variant of the alternative Data–Estimation method (see Appendix˜C for details). This gives an approach that is similar to those of Hu et al, (2024); Piriyakulkij et al, (2023); Wang et al, (2025); Chan et al, (2025), but which does not require us to assume a single fixed set of possible hypotheses throughout the sequential BED process. We evaluate this alternative approach on the Celebrities dataset using Qwen2.5-72B as both the questioner and answerer. Results are shown in Figure 4. Consistent with our analysis in Section˜4, BED-LLM substantially outperforms the Data–Estimation alternative. Data–Estimation still outperforms the Naive QA baseline, indicating that even this pairing benefits from explicit EIG–driven question selection over purely heuristic question generation, but interestingly does worse than Entropy, emphasizing the importance of estimating our uncertainty in the $y$ space instead of $\theta$ space. These findings reinforce our claim that the choice of joint–model factorization is a critical design decision for sequential BED in the LLM setting.
6.2 Active Preference Elicitation
In our 20–Questions problems, the generated hypotheses were all single, concrete entities (barring any failures of the LLMs) and most reasonable questions had clear–cut answers (e.g. “is Beyoncé a singer?”). In contrast, many problems, especially those related to eliciting information from a user, involve more abstract targets and/or more subjective and less predictable answers. For example, if we are learning about the characteristics or preferences of a user, then it may be difficult to explicitly define a concrete closed set of possible $\theta$, while it is also quite challenging for the LLM to accurately predict user responses and develop appropriate uncertainty estimates.
To study such a scenario, we evaluate BED-LLM on inferring users’ film preferences. Here the target we are trying to gather information about is somewhat abstract, and we have some flexibility in how we define $\theta$ in our joint model. Our chosen setup is to define $\theta$ to be a natural language paragraph describing a user’s taste in films; we refer to this as a persona. To emulate a human user, our answerer is set up to be an LLM prompted to answer questions as if it were a person with a given persona $\theta^{*}$.
To generate a dataset of ground-truth personas, we take a set of $200$ real user ratings from the MovieLens-100K dataset (Harper & Konstan,, 2015), then use an “oracle” LLM (namely, OpenAI’s o3 model) to produce a natural language persona that is consistent with each distinct set of ratings. As was the case for the 20–Questions problems, this dataset is never provided to the questioner and the set of allowed $\theta$ is not constrained.
Because we need the LLM to be able to meaningfully capture uncertainty in the space of responses as discussed in Section˜4.2, we restrict questions to be multiple-choice. Specifically, the questioner is tasked with producing a question along with five possible responses A/B/C/D/E. We then define each $x_{t}$ to be the question coupled with the possible responses, and each $y_{t}$ to be one of the letters A/B/C/D/E to provide a restricted set of tokens over which we can measure entropy. Option E is further constrained to always be “none of the above” so that the answerer is not committed to choosing one of the directly generated choices if none are suitable.
For BED-LLM and Entropy, we compare $M=8$ candidate questions at each turn and we aim to generate at least $N=5$ candidate hypotheses. We use the “unconstrained generation” approach of candidate question generation (c.f. Section˜3.1) as the user profiles can be quite diffuse and we are only generating a small number of possible hypotheses that can be quite easy to split.
We note that data–estimation setup is not at all viable for this problem because the large number of tokens and varying dimensionality of each $\theta$ sample mean that $\mathrm{H}[p_{\mathrm{LLM}}(\theta;[h_{t},x_{t+1},y_{t+1}])]$ is not only infeasible to estimate, but also is not a meaningful measure of uncertainty.
(a) Qwen2.5-72B
(b) GPT-4o-mini
(c) Llama-3.3-70B
(d) GPT-4o
Figure 5: Results for the active preference elicitation with the same LLM used for the questioner and answerer. Shown is the average ratings of 10 recommendations for each of 200 separate simulated users.
Evaluation
To again allow tracking of progress through the experiment, after each turn of the interaction $t$, the questioner generates ten film recommendations, conditioned on $h_{t}$. These recommendations are checked for consistency with prior questions and answers; if any are judged inconsistent then they are removed and additional recommendations are generated. The quality of the film recommendations is then assessed using an “LLM-as-judge” protocol (Zhu et al,, 2025; Trivedi et al,, 2024). Namely, the answerer evaluates each of the 10 films recommended by the questioner, conditioned on the hidden target user profile $\theta^{*}$. It scores each film on a scale of 1 to 5 (in 0.5 increments), based on how well the recommendation aligns with $\theta^{*}$ — this score is output together with a brief justification to increase reliability. We report the mean rating and standard error across 200 users, over 5 question–answer turns.
Results
In Figure˜5, we show the results using the same model for both questioner and answerer. For all models, BED-LLM establishes an advantage over other baselines by turn $t=3$, although the differences in performance between methods are smaller than for the 20–Questions game. On average, Entropy performs similarly to Naive QA.
(a) Q: Qwen2.5-72B. A: GPT-4o-mini.
(b) Q: GPT-4o-mini. A: Qwen2.5-72B.
Figure 6: Results for Active Preference Elicitation with a different models: questioner Q and answerer A.
Different questioner and answerer
Like in the 20–Questions setting, we now also evaluate performance on Active Preference Elicitation in the misspecified model scenario by using a different LLM for the questioner and answerer. Figure˜6 presents the results. While BED-LLM and Entropy seem to be quite robust to the model misspecification, it is now Naive QA that seems to be brittle, exhibiting a drop in performance when using Qwen2.5-72B questioner and GPT-4o-mini answerer.
7 Conclusion
In this work, we have shown how to effectively apply the framework of sequential Bayesian experimental design (BED) to the problem of interactive information gathering with LLMs. Specifically, we have introduced BED-LLM, which provides a specific, information–theoretic, sequential BED approach that makes a variety of carefully justified design choices in the joint-model factorization, belief updating, and candidate question generation. Particularly central to BED-LLM is the prior–likelihood pairing with filtering of hypotheses for consistency with the history. BED-LLM is notably the first work that uses both this prior–likelihood pairing without resorting to using marginal predictive entropy in lieu of the expected information gain (EIG). Together, these innovations lead to substantial performance improvements compared to previous approaches. The results thus confirm that principled EIG–driven strategies can yield substantial gains for interactive, multi–turn, information gathering problems.
Appendix A Additional discussions on design choices
A.1 An alternative view on the faithfulness of conditional distributions
Another way of viewing the distinction between the prior–likelihood and data–estimation constructions is in which of the EIG forms, Equation˜2 or Equation˜3, we center our reasoning. For a given joint model, the two are, of course, mathematically equivalent. However, they give us different ways of thinking about what it means to maximize the EIG: reducing entropy in $\theta$ from seeing $y$, or reducing entropy in $y$ from seeing $\theta$. This, in turn, gives us a way to reason about how appropriate our joint model is. When we choose to use one of $p\left(y_{t+1};\left[\theta,x_{t+1}\right]\right)$ or $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$, we are centering our reasoning around the entropy of this quantity making sense, while allowing the other entropy in the other form to be implicitly defined from the resulting joint distribution; because the two forms are equivalent, we know that if our explicit form is suitable/unsuitable, the implicit form will be as well. If, for example, we directly fix the form of $p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$ using our LLM’s predictive distribution, we are also directly relying on its expected entropy being a meaningful measure of design quality. If $\theta$ is high–dimensional and predominantly free–form, the resulting entropy produced by the LLM is unlikely to be meaningful and using the data–estimation pairing is unlikely to produce an effective strategy. However, if $y$ is instead quite constrained, the LLM can produce a meaningful entropy over it, and choosing a model based on the prior–likelihood pairing is likely to implicitly define a meaningful distribution, and thus entropy, on $\theta$. Conversely, if $\theta$ is constrained and $y$ is free form, the opposite will hold instead.
A.2 Choice of $\theta$
An important corollary of this reasoning is that it can be important to be careful in our choice of exactly what we take $\theta$ to be, especially if we are using the data–estimation formulation. In particular, it is essential for entropy in the space of $\theta$ to form a meaningful notion of uncertainty, even if this entropy is not being measured through the LLM’s predictive distribution of $\theta$ directly. Thus, while $\theta$ inherently represents what we are trying to learn about and should always be set up as such, if there is flexibility in how exactly we formulate it, we should be careful to choose a form that yields an appropriate uncertainty measure. For example, if the LLM is trying to clarify what code a user wishes it to generate, we could either choose $\theta$ to be the code itself or, following (Neiswanger et al,, 2021; Bickford Smith et al,, 2023), the output the code produces. Here the entropy over code outputs induced by our distribution on code is likely to be a much better measure of uncertainty than the entropy of the raw code itself, given that there are multiple ways one can code the same operation.
A.3 Expected uncertainty reduction in the sequential setting
In Section˜4.4, we looked at whether the prior–likelihood pairing or the data–estimation pairing is better aligned with how we update our beliefs in practice. The picture here can get a somewhat more complicated once we move into the sequential BED setting. Here, our ultimate aim is actually to minimize $\mathrm{H}[p(\theta;h_{T})]$ at some final future horizon $T$. Now, we only care about intermediary belief states $p(\theta;h_{t})$ through their aid in future decision making toward the goal of minimizing the final entropy. Thus, even if we are working with in-context updates, it might be the case that $p(\theta;h_{t})$ only starts to produce a meaningful entropy once we have seen enough data to sufficiently narrow down the possibilities on $\theta$. The optimal behavior in such settings would be to learn a policy that directly targets this final belief state instead of sequentially targeting the incremental EIGs. However, this will typically not be computationally feasible in practice and we instead need to resort to a myopic decision-making strategy. It might thus still be better to use the prior–likelihood formulation in such myopic decision making settings, even if we are sequentially updating our beliefs on $\theta$ through in-context updates, if this allows us to better guide the sequential decisions towards our final objective. The coherence of Bayesian updating means that the converse is unlikely to be true, so this provides further evidence towards using the prior–likelihood formulation.
Appendix B Generating candidate hypotheses for BED-LLM
A fundamental challenge for BED-LLM and the same method with the Entropy baseline is generating a sufficiently diverse set of candidate hypotheses from the LLM prior. As illustrated in Figure˜8, the raw distribution $p_{\mathrm{LLM}}(\theta)$ is highly over–collapsed, often concentrating mass on only a few high–likelihood hypotheses. Thus, it is not practical to directly use the LLM’s distribution as a prior $p(\theta)$. Instead, we jointly sample candidate $\theta$ and assume a uniform distribution over them. We can view this as sampling thetas from a mixture distribution.
Instead, we employ a joint sampling strategy, in which the LLM is prompted to generate a list of $N$ hypotheses in a single rollout, which corresponds to drawing from the autoregressive list distribution
$p_{\mathrm{LLM}}(\theta^{(1)}_{t},\ldots,\theta^{(N)}_{t};h_{t})\;=\;\prod_{n=1}^{N}p_{\mathrm{LLM}}(\theta_{t}^{(n)};[\theta_{t}^{(1:n-1)},h_{t}])$
Each such joint sample is then filtered for history consistency using $p_{\mathrm{LLM}}(h_{t})$. This may be repeated. The surviving hypotheses from the current rollout are then merged with those retained from previous steps to form a cumulative candidate set $\Theta$. While one could in principle reweight candidates using importance sampling, in practice we choose to not rely on the model’s internal probabilities. Instead approximate the prior as a uniform distribution over this union
$p_{f}(\theta;h_{t})\approx\frac{1}{\Theta}\sum_{\theta\in\Theta}\delta_{\theta}.$
To further improve coverage and diversity, we incorporate several refinements in the generation of $p_{f}(\theta;h_{t})$. Due to the differences between different LLMs, the ways this is encouraged may differ, but we find in general that the following are effective:
  • Repeat sampling and filtering: The sampling, filtering, and merging steps may be repeated to encourage diversity and ensure filtering correctness.
  • Prompts: Prompts are designed to elicit stratified hypotheses by encouraging the LLM to consider different semantic features (e.g. age groups, genres, or categories) and implicitly diversify across them.
  • Context ordering: During generation, we reverse the order of the question–answer pairs in $h_{t}$ to place the most recent question at the top of the context window (while retaining earlier exchanges), ensuring that specific constraints are prioritized and mitigating context drift.
Finally, we note that different LLMs respond differently to strategies aiming to increase diversity: some benefit more from a higher temperature while others benefit from more repetitions of the sampling–filtering cycle.
Appendix C Data–estimation method
Our Data–Estimation method is based on a model derived from a data–estimation pairing (Section˜3).
C.1 EIG Estimation
Suppose our model is given by $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p\left(y_{t+1};[h_{t},x_{t+1}]\right)p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)$}. Here, it will clearly be beneficial to directly use Equation˜2 for estimating the EIG, as here we directly have access to all the required terms, other than $p(\theta)$ which can be simply ignored as it is not a function of $x$ so does not affect optimization of the question. If the possible values for $y$ are enumerable and we can evaluate $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$ in closed–form, we can directly calculate the exact EIG (up to a constant) without requiring any estimation at all:
$\text{EIG}_{\theta}(x)-\mathrm{Const}=-\sum_{y}p\left(y_{t+1};[h_{t},x_{t+1}]\right)\mathrm{H}\left[p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)\right],$ (8)
where the entropy $\mathrm{H}\left[p\left(\theta;[h_{t},x_{t+1},y_{t+1}]\right)\right]$ can be evaluated directly from the logits of the LLM, or if these are not available, estimated by sampling. If we cannot enumerate $y$ or evaluate $p\left(y_{t+1};[h_{t},x_{t+1}]\right)$, we can simply instead resort to Monte Carlo and use the estimator:
$\text{EIG}_{\theta}(x)-\mathrm{Const}\approx-\frac{1}{N}\sum_{n=1}^{N}\mathrm{H}\left[p(\theta;[c,x,y_{n}])\right]\quad\mathrm{where}\quad y_{n}\sim p\left(y_{t+1};[h_{t},x_{t+1}]\right).$ (9)
We provide an overview of how to do this in the LLM setting in Algorithm˜1.
Algorithm 1 Data–Estimation Selection at Turn $t$
1:History $h_{t}$; candidate questions $\mathcal{X}_{\text{cand}}^{t}$; answer sets $\{\mathcal{Y}(x)\}$
2:Selected question $x_{t}^{\star}$
3:for each $x\in\mathcal{X}_{\text{cand}}^{t}$ do
4:  Obtain predictive answer distribution $p_{\text{LLM}}(y;x,h_{t})$ for all $y\in\mathcal{Y}(x)$
5:  for each $y\in\mathcal{Y}(x)$ do
6:   Construct $\Theta_{t,y}\sim p_{f}(\theta;h_{t}\cup(x,y))$
7:   Compute entropy $H_{y}\leftarrow\mathrm{H}\!\left[p_{\mathrm{LLM}}(\theta;h_{t},x,y)\right]$
8:  end for
9:  Compute $\text{EIG}(x)\leftarrow-\sum_{y\in\mathcal{Y}(x)}p_{\text{LLM}}(y;x,h_{t})\;H_{y}$
10:end for
11:Select $x_{t}^{\star}\leftarrow\arg\max_{x\in\mathcal{X}_{\text{cand}}^{t}}\text{EIG}(x)$
12:return $x_{t}^{\star}$
C.2 Generating candidate hypotheses
To generate candidate values of $\theta$ for the data–estimation method, we used the prompt in Figure˜7. Figure˜8 shows an example of the distribution of samples obtained following two rounds in the 20–questions game. Note that the samples are highly concentrated on just a handful of answers. This lack of diversity shows that the model’s belief distribution is far more concentrated relative to the variability over valid hypotheses in the ground–truth task distribution, which negatively impacts the performance of the data–estimation method.
Return only the full name of one randomly selected famous person (living or deceased) consistent with the questions and answers above.
To increase randomness:
1. Internally brainstorm a pool of diverse and representative individuals.
2. Avoid defaulting to the most globally ubiquitous celebrities or famous figures.
Output rules:
- Output ONLY the person’s full name (with spaces, capitalization and accents), nothing else.
- No extra words, explanations, numbering, or punctuation beyond what’s in the name itself (hyphens/apostrophes allowed if part of the name).
Figure 7: Prompt for generating hypotheses (and evaluating their probability) for the data–estimation method.
Is this person known for their contributions to science?
No.
Is this person known for their contributions to the arts?
Yes.
{"Vincent van Gogh": 93,
"Georgia O Keeffe": 10,
"Frida Kahlo": 37,
"Salvador Dali": 44,
"August Wilson": 8,
"Auguste Rodin": 8}
Figure 8: An example of the sample distribution generated using the prompt in Figure˜7, conditioned on the two question/answer pairs at the top of this figure. At this stage of the game, we independently sample 200 hypotheses from the LLM and record their frequencies. Note that the distribution exhibits strong mode collapse, with most of the mass highly concentrated on just a few answers, which negatively impacts the performance of the data–estimation method. This summary is for diagnostic purposes: Algorithm˜1 operates on the probabilities of individual samples and never instantiates such an aggregated summary.
Appendix D Experimental details
D.1 Generation of the Animals Dataset
To create the Animals dataset, we prompted OpenAI o3 (OpenAI,, 2025, o3-2025-04-16) to generate a list of animals, using the prompt in Figure˜9. The resulting list is shown in Figure˜10. Alternative names (after $|$ character) were manually added.
You are a zoologist. Please generate a list of 100 living animal species with very high taxonomic diversity, including diversity in phyla, classes, orders, and families. Present each animal on a different line.
Figure 9: Prompt for Animals dataset generation.
African elephant
Bengal tiger
Bald eagle
Blue whale
Red kangaroo
Giant panda
Snow leopard
Green sea turtle
American alligator
Bottlenose dolphin
Emperor penguin
Great white shark
Golden poison frog | Golden poison dart frog
Honey bee
Monarch butterfly
Okapi
Chimpanzee
Arctic fox
Komodo dragon
Giraffe
Cheetah
Hammerhead shark
Axolotl
Orca
Puffin
Red panda
Platypus
Rhinoceros beetle
Tasmanian devil
Wombat
Sloth
Blue-ringed octopus | Blue ringed octopus
Manatee
Narwhal
Sea otter
Coral snake
King cobra
Harpy eagle
Lemur
Koala
Aye-aye | Ayeaye
Snowy owl
Elk
Wolverine
Caracal
Cassowary
Quokka
Pangolin
Saiga antelope
Galápagos tortoise | Galapagos tortoise
Sumatran orangutan
Red-eyed tree frog | Redeyed tree frog
European badger
Moose
African grey parrot
Scarlet macaw
Black mamba
Albatross
Humpback whale
Dugong
Anaconda
Kookaburra
Coyote
Brown bear
Golden jackal
Capybara
Ibex
Japanese macaque
Kiwi
Leafcutter ant
Mantis shrimp
Ocelot
Peregrine falcon
Quetzal
Raccoon
Sand cat
Tarantula
Uakari
Vicuña
Wildebeest
Rock hyrax | dassie
Yak
Zebra
Blue dragon nudibranch | Blue dragon sea slug
Chinchilla
Dhole
Electric eel
Flying fox
Gharial
Horseshoe crab
Indigo bunting
Jerboa
Kakapo
Lionfish
Markhor
Nautilus
Olive baboon
Pika
Quoll
Rosy boa
Figure 10: Animals dataset (generated using OpenAI o3, with manual curation)
D.2 LLM models details
The open-weights LLM models used in Section˜6 were obtained from HuggingFace (HF) (Wolf et al,, 2020) and their repositories are: mistralai/Mistral-Large-Instruct-2411, Qwen/Qwen2.5-72B-Instruct, meta-llama/Llama-3.3-70B-Instruct. Since all models are instruction-finetuned, we used the latest chat templates provided as part of the HF implementation. All of the experiments using mistralai/Mistral-Large-Instruct-2411 were conducted by Oxford researchers on Oxford compute infrastructure.
The OpenAI API models used were: gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06.
Appendix E Additional Personalization Experiments
E.1 EIG estimates of questions chosen by different questioners
Figure 11 demonstrates the progression of EIG estimates throughout the turns for BED-LLM and Naive QA. We see that at turn $t=1$, the questions chosen by both methods have similar EIG estimates, which explains the similar performance of BED-LLM and Naive QA at turn $t=1$.
Figure 11: The average information gains for each strategy at each turn for the Qwen2.5-72B vs Qwen2.5-72B example. We see that the initial questions have similar information gains, which explains why the performance of the Naive QA method outperforms EIG at turn 1. Afterwards, the EIG information gains are much larger while the Baseline information gains decrease sharply.